OpenStudy (anonymous):

check my answer plz? State whether the given measurements determine zero, one, or two triangles. A = 36°, a = 3, b = 9 I got zero

OpenStudy (mathstudent55):

Correct.

OpenStudy (anonymous):

Thank you! Can you also check two more questions for me?

OpenStudy (anonymous):

1. Determine whether the vectors u and v are parallel, orthogonal, or neither. u = <1, -2>, v = <-4, 8> I got Neither. 2. Evaluate the expression. r = <9, -7, -1>, v = <2, 2, -2>, w = <-5, -2, 6> v ⋅ w I got <-10, -4, -12>

OpenStudy (mathstudent55):

As you can see, a = 3 is too small to make a triangle. 

OpenStudy (anonymous):

Yes, a was 3 and b was 9 as indicated int he question

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):

Slope of <1, 2> = 2/1 = 1 Slope of <-4, 8> = -8/4 = -2 Since 1 and -2 are not negative reciprocals, the vectors are not perpendicular. Since 1 and -2 are not equal, the vectors are not parallel.

OpenStudy (anonymous):

Got it

OpenStudy (anonymous):

And my last one is okay as well?

OpenStudy (anonymous):

@mathstudent55

OpenStudy (mathstudent55):

The dot product of two vectors is a scalar, just a number.

OpenStudy (mathstudent55):

v ⋅ w = <9, -7, -1>⋅ <-5, -2, 6> = 9(-5) + (-7)(-2) + (-1)(6) = -45 + 14 - 6 = -37